# Could Small Language Models Serve as Recommenders?     Towards Data-centric Cold-start Recommendations

### 16 Feb 2024 v4

@@ Abstract.
Recommendation systems help users find information that matches their interests based on their historical behaviors. However, generating personalized recommendations becomes challenging in the absence of historical user-item interactions, a practical problem for startups known as the system cold-start recommendation. Current research tackles user or item cold-start scenarios but lacks solutions for system cold-start. To tackle the problem, we initially propose PromptRec, a simple but effective approach based on in-context learning of language models, where we transform the recommendation task into the sentiment analysis task on natural language containing user and item profiles. However, this naïve strategy heavily relied on the strong in-context learning ability emerged from large language models, which could suffer from significant latency for online recommendations. To fill this gap, we present a theoretical framework to formalize the connection between in-context recommendation and language modeling. Based on it, we propose to enhance small language models with a data-centric pipeline, which consists of: (1) constructing a refined corpus for model pre-training; (2) constructing a decomposed prompt template via prompt pre-training. They correspond to the development of training data and inference data, respectively. To evaluate our proposed method, we introduce a cold-start recommendation benchmark, and the results demonstrate that the enhanced small language models can achieve comparable cold-start recommendation performance to that of large models with only around $17\%$ of their inference time. To the best of our knowledge, this is the first study to tackle the system cold-start recommendation problem. We believe our findings will provide valuable insights for future works. The benchmark and implementations are available at https://github.com/JacksonWuxs/PromptRec.
In-context learning; cold-start recommendation; language models; data-centric AI.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singapore.booktitle: Proceedings of the ACM Web Conference 2024 (WWW ’24), May 13–17, 2024, Singapore, Singaporeprice: 15.00isbn: 979-8-4007-0171-9/24/05doi: 10.1145/XXXXXX.XXXXXXjournalyear: 2024copyright: rightsretainedconference: Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Proceedings of the ACM Web Conference 2024 (WWW ’24), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589334.3645494isbn: 979-8-4007-0171-9/24/05ccs: Information systems Personalization
@@ 1. Introduction
One potential direction to solve the system cold-start recommendation problem is leveraging in-context learning with large language models (LLMs) (Vaswani et al., 2017; Brown et al., 2020; Liu et al., 2021), demonstrated by LLMs could quickly adapt to a new task (recommendation in our study) without training them on a task dataset. Intuitively, the relation between user interests and item properties has been implicitly expressed as a natural language in public corpora. Thus, it could be captured by LLMs and used for cold-start recommendations. However, in-context learning is an emergent ability of large language models, which usually starts from hundreds of millions to billions of parameters. Therefore, adopting these methods for online recommendation is impractical due to their slow and costly inference. This naturally raises a question: could small language models be in-context recommenders for system cold-start recommendation?
@@
To analyze the reason behind this failure, we propose a theoretical framework to formalize the mechanism of in-context recommendation with PromptRec under the Hidden Markov Model (HMM) assumption (Xie et al., 2021). Under the assumption, a language model first infers the “concept” (sentiment polarity in this study) based on the input prompt (user-item profiles and a template in this study), and then makes recommendations conditioned on both the inferred concept and the input prompt. Our analysis demonstrates that language models perform in-context recommendations by estimating the likelihood of the sentiment words conditioned on the user-item context and the prior probability to different pre-trained concepts. Small language models fail to provide precise estimations of these two factors due to their limited parameters.
With this finding, we enhance the small language models’ recommendation performance with improved estimation of in-context prediction probabilities via a data-centric pipeline, including two major steps: model pre-training and prompt template pre-training. Specifically, if we could (1) pre-train the language model on a corpus related to the recommendation scenario, and (2) pre-train the prompt template with interactions from other recommendation domains, the cold-start recommendation performance of small language models could be enhanced. Note that the training data belongs to distributions different from the target recommendation domain, so it does not violate the system cold-start setting. However, we face two challenges due to the lack of user-item interactions in the cold-start scenario: (1) how to find a corpus to pre-train small models for recommendation tasks? (2) how to make the prompt template generalizable across different domains? For the first challenge, we propose refining a general corpus by maximizing mutual information between the general documents and the synthetic user-item interactions in the cold-start scenario. For the second challenge, we propose to decompose prompt templates into the “task” and “domain” prompts, where the former is transferable across different recommendation scenarios. We verify these two methods and show that the enhanced small language models achieve comparable performance with large models in just around $17\%$ of their inference times. Remarkably, enhanced BERT-mini with $11.3$M parameters ($\approx$3of BERT-large) achieve performance comparable with BERT-large in terms of in-context recommendations under the system cold-start setting. We summarize contributions as follows:
@@ 3. Methodology
We introduce the details of the proposed framework. In Sec. 3.1, we present PromptRec, an in-context learning approach to tackle the system cold-start problem. In Sec. 3.2, we provide a theoretical framework to build the connection between in-context recommendation and language model. Furthermore, we develop a data-centric pipeline to enhance the in-context recommendation of (small) language models, with a focus on training corpus refinement in Sec. 3.3, and inference-time prompt design in Sec. 3.4.
@@ 4.3.1. Experiment Designs
Details of transferable prompt pre-training (TPPT). For each dataset in the benchmark, we treat it as a single cold-start scenario, so we use the other two datasets to pre-train the task prompt $\mathbf{S}_{t}$. We adopt a continuous task prompt, represented as trainable vectors aligned with the pre-trained model’s dimension. Our training utilizes the AdamW optimizer with a learning rate of $1e^{-4}$ and a batch size of 64 for 50K steps. We validate the performance on the validation set at the end of each epoch. If the performance decreases consecutively twice, the learning rate will be reduced by a factor $0.5$. The early stop strategy is triggered by three 3 times learning rate reduction. Domain prompts $\mathcal{S}_{d}$ derive from each dataset’s top TF-IDF (Salton and Buckley, 1988) scored keywords. Every task has a 50-word task prompt, while domain prompts vary: 5, 10, 50, 100, and 200 words. Experiments for each configuration use five random seeds, with results reported for the best-performing average setup.
@@ 4.3.2. Experiment Results
Pre-training transferable prompts help small language models for better cold-start recommendation performance.
In Table 4, the TPPT strategy markedly enhances PromptRec’s cold-start recommendation performance across all situations. Specifically, BERT-small on the coupon dataset attains a GAUC of 64.57, surpassing the highest GAUC of 63.77 achieved by BERT-large among all candidate models. As the same observation on RCMP, TPPT facilitates in-context recommendations in all five settings where previously the language models were suboptimal. Furthermore, TPPT consistently outperforms RCMP in cold-start scenarios. These results indicate that while small language models possess the potential for in-context recommendation in cold-start situations, they require optimal prompts for activation.
Improved small language models make effective cold-start recommendations with an efficient inference speed.
@@ 5.1. Cold-start Recommendation
The phrase “cold-start“ describes the situation when the recommender knows nothing about its serving objects. There are several cases of cold-start recommendations.
@@ 5.2. In-context Recommendation
Researchers leverage in-context learning to develop recommenders for several purposes, including interpretability (Li et al., 2022; Gao et al., 2023), multi-tasks learning (Geng et al., 2022; Zhang et al., 2023; Liu et al., 2023a), sequential recommendation (Zhang et al., 2021; Sileo et al., 2022; Cui et al., 2022; Hou et al., 2023; Petrov and Macdonald, 2023; Wang et al., 2023), and conversational recommendation (Wang et al., 2022; Gao et al., 2023). The intuitions behind these studies are two folds: utilizing language models as a knowledge base to enhance recommendation (Zhang et al., 2021; Sileo et al., 2022; Cui et al., 2022; Zhang et al., 2023; Liu et al., 2023a); using language models as a bridge to realize the interaction between humans and systems with natural language (Wang et al., 2022; Li et al., 2022; Geng et al., 2022; Hou et al., 2023; Petrov and Macdonald, 2023; Wang et al., 2023; Gao et al., 2023). These studies inspire us to push in-context learning to a more challenging scenario, where the user’s shopping history is not available. Moreover, in contrast with their focus on large language models, our study demonstrates that small language models could also provide recommendations under the cold-start setting.
@@ 6. Conclusions
This paper studies the system cold-start recommendation problem, including providing a formal definition and the first benchmark. We propose PromptRec for this problem and show that large language models can make personalized recommendations without any training samples. In addition, we provide a mathematical framework to study the behavior of language models to make in-context recommendations. We also investigate two methods to improve small language models by leveraging the datasets that are out-of-domains of the cold-start scenario. Our results demonstrate that the small language models could make personalized recommendations with their enhanced in-context ability. We encourage future research to explore the system cold-start setting for more recommendation tasks and hope they could be deployed in real-world businesses.

### 04 Mar 2024 v5

@@ Abstract.
Recommendation systems help users find matched items based on their previous behaviors. Personalized recommendation becomes challenging in the absence of historical user-item interactions, a practical problem for startups known as the system cold-start recommendation. While existing research addresses cold-start issues for either users or items, we still lack solutions for system cold-start scenarios. To tackle the problem, we propose PromptRec, a simple but effective approach based on in-context learning of language models, where we transform the recommendation task into the sentiment analysis task on natural language containing user and item profiles. However, this naïve approach heavily relies on the strong in-context learning ability emerged from large language models, which could suffer from significant latency for online recommendations. To solve the challenge, we propose to enhance small language models for recommender systems with a data-centric pipeline, which consists of: (1) constructing a refined corpus for model pre-training; (2) constructing a decomposed prompt template via prompt pre-training. They correspond to the development of training data and inference data, respectively. The pipeline is supported by a theoretical framework that formalizes the connection between in-context recommendation and language modeling. To evaluate our approach, we introduce a cold-start recommendation benchmark, and the results demonstrate that the enhanced small language models can achieve comparable cold-start recommendation performance to that of large models with only $17\%$ of the inference time. To the best of our knowledge, this is the first study to tackle the system cold-start recommendation problem. We believe our findings will provide valuable insights for future works. The benchmark and implementations are available at https://github.com/JacksonWuxs/PromptRec.
In-context learning; cold-start recommendation; data-centric AI; large language models.
@@ 1. Introduction
One potential direction to solve the system cold-start recommendation problem is leveraging in-context learning with large language models (LLMs) (Vaswani et al., 2017; Brown et al., 2020; Liu et al., 2021), as LLMs can easily adapt to a new task (recommendation in our study) without training them on a task dataset. Intuitively, the relation between user interests and item properties has been implicitly expressed as a natural language in public corpora. Thus, it could be captured by LLMs and used for cold-start recommendations. However, in-context learning is an emergent ability of large language models, which usually starts from hundreds of millions to billions of parameters. Therefore, adopting these methods for online recommendation is impractical due to their slow and costly inference. This naturally raises a question: Could small language models be in-context recommenders for system cold-start recommendation?
@@
To understand the reason, we propose a theoretical framework to formalize the mechanism of in-context recommendation with PromptRec under the Hidden Markov Model (HMM) assumption (Xie et al., 2021). Under the assumption, a language model first infers the “concept” (sentiment polarity) based on the input prompt (user-item profiles and a template), and then makes recommendations conditioned on both the inferred concept and the input. Our analysis demonstrates that language models perform in-context recommendations by estimating the likelihood of the sentiment words conditioned on the user-item context and the prior probability to different pre-trained concepts. Small language models fail to provide precise estimations of these two factors due to their limited parameters.
With this finding, we enhance the small language models’ recommendation performance with improved estimation of in-context prediction probabilities via a data-centric pipeline, including two major steps: model pre-training and prompt template pre-training. Specifically, if we could (1) pre-train the language model on a corpus related to the recommendation scenario, and (2) pre-train the prompt template with interactions from other recommendation domains, the cold-start recommendation performance of small language models could be enhanced. Note that the training data belongs to different distributions from the target recommendation domain, so it does not violate the system cold-start setting. However, we face two challenges due to the lack of user-item interactions in the cold-start scenario: (1) how to find a corpus to pre-train small models for recommendation tasks? (2) how to make the prompt template generalizable across different domains? For the first challenge, we propose refining a general corpus by maximizing mutual information between the general documents and the synthetic user-item interactions in the cold-start scenario. For the second challenge, we propose to decompose prompt templates into the “task” and “domain” prompts, where the former is transferable across different recommendation scenarios. We verify these two methods and show that the enhanced small language models achieve comparable performance with large models in just around $17\%$ of their inference times. Remarkably, enhanced BERT-mini with $11.3$M parameters ($\approx$3of BERT-large) achieve performance comparable with BERT-large in terms of in-context recommendations under the system cold-start setting. We summarize contributions as follows:
@@ 3. Methodology
We now introduce the details of the proposed approach. In Sec. 3.1, we present PromptRec, an in-context learning method to tackle the system cold-start problem. In Sec. 3.2, we provide a theoretical framework to build the connection between in-context recommendation and language model. Furthermore, we develop a data-centric pipeline to enhance the in-context recommendation of (small) language models, with a focus on training corpus refinement in Sec. 3.3, and inference-time prompt design in Sec. 3.4.
@@ 4.3.1. Experiment Designs
Details of transferable prompt pre-training (TPPT). For each dataset in the benchmark, we treat it as a single cold-start scenario, so we use the other two datasets to pre-train the task prompt $\mathbf{S}_{t}$. We adopt a continuous task prompt, represented as trainable vectors aligned with the pre-trained model’s dimension. Our training utilizes the AdamW optimizer with a learning rate of $1e^{-4}$ and a batch size of 64 for 50K steps. We validate the performance on the validation set at the end of each epoch. If the performance decreases consecutively twice, the learning rate will be reduced by a factor $0.5$. The early stop strategy is triggered by three 3 times learning rate reduction. Domain prompts $\mathcal{S}_{d}$ derive from each dataset’s top TF-IDF (Salton and Buckley, 1988) scored keywords. The length of the task prompt is chosen as 10 or 50, while that of domain prompts varies: 5, 10, 50, 100, and 200. Experiments for each configuration use five random seeds, with results reported for the best-performing average setup.
@@ 4.3.2. Experiment Results
Pre-training transferable prompts help small language models for better cold-start recommendation.
In Table 4, the TPPT strategy enhances PromptRec’s cold-start recommendation performance in most situations. Specifically, BERT-small on the coupon dataset attains a GAUC of 64.57, surpassing the highest GAUC of 63.77 achieved by BERT-large among all candidate models. Also, TPPT facilitates in-context recommendations in all five settings where previously the language models were suboptimal. Furthermore, TPPT consistently outperforms RCMP in cold-start scenarios. These results indicate that while small language models possess the potential for in-context recommendation in cold-start situations, they require optimal prompts for activation.
Improved small language models make effective recommendations with a significantly faster inference speed.
@@ 5.1. Cold-start Recommendation
The phrase “cold-start” describes the situation when the recommender knows nothing about its serving objects. There are several cases of cold-start recommendations.
@@ 5.2. In-context Recommendation
Researchers leverage in-context learning to develop recommenders for several purposes, including interpretability (Li et al., 2022; Gao et al., 2023), multi-tasks learning (Geng et al., 2022; Zhang et al., 2023; Liu et al., 2023a), sequential recommendation (Zhang et al., 2021; Sileo et al., 2022; Cui et al., 2022; Hou et al., 2023; Petrov and Macdonald, 2023; Wang et al., 2023), and conversational recommendation (Wang et al., 2022; Gao et al., 2023). The intuitions behind these studies are two folds: utilizing language models as a knowledge base to enhance recommendation (Zhang et al., 2021; Sileo et al., 2022; Cui et al., 2022; Zhang et al., 2023; Liu et al., 2023a); using language models as a bridge to realize the interaction between humans and systems with natural language (Wang et al., 2022; Li et al., 2022; Geng et al., 2022; Hou et al., 2023; Petrov and Macdonald, 2023; Wang et al., 2023; Gao et al., 2023). In contrast with their focus on large language models, our study demonstrates that small language models could also provide recommendations under the cold-start setting.
@@ 6. Conclusions
This paper studies the system cold-start recommendation problem and provides the first benchmark. We propose PromptRec for this problem and show that large language models can make personalized recommendations without any training samples. In addition, we provide a mathematical framework to study the behavior of language models to make in-context recommendations. We also investigate two methods to improve small language models by leveraging the datasets that are out-of-domains of the cold-start scenario. Our results demonstrate that the small language models could make personalized recommendations with their enhanced ability. We encourage future research to explore the system cold-start setting for more recommendation tasks and hope they could be deployed in real-world businesses.